OpenStudy (anonymous):

Answer please !! The length of a rectangular prism is twice the width.The height of the prism is equals to width.Find the dimensions of the prism,given that the volume is 54 cubic inches.

OpenStudy (amistre64):

"Answer please" might be a bit misleading; either that or you dont understand the nature of this site ....

OpenStudy (amistre64):

drawing it out might prove useful

OpenStudy (amistre64):



OpenStudy (amistre64):

we are given the h = w

OpenStudy (amistre64):

do you know how to find the volume of a rectangular prism?

OpenStudy (anonymous):

V=LWH

OpenStudy (amistre64):

good, and since w=h lets say: V = LW^2

OpenStudy (amistre64):

we know V = 54 soo 54 = LW^2 and it says L = 2*W

OpenStudy (amistre64):

54 = 2WWW = 2W^3 then right?

OpenStudy (amistre64):

solve for W and you will have the means to get the other measures W = W H = W L = 2W

OpenStudy (anonymous):

ah .. i get it !

OpenStudy (anonymous):

then ?